A relatively small impact crater 20 kilometers in diameter could be made by a comet 2 kilometers in diameter traveling at 25 km/s (2.5×104 m/s ).

a) Assume that the comet has a total mass of 4.2×1012 kilograms. What is its total kinetic energy? (Hint: The kinetic energy is equal to 12mv2, where m is the comet's mass and v is its speed. If you use mass in kilograms and velocity in m/s, the answer for kinetic energy will have units of joules.) Express your answer using two significant figures.

b)Convert your answer from part A to an equivalent in megatons of TNT, the unit used for nuclear bombs. Comment on the degree of devastation the impact of such a comet could cause if it struck a populated region on Earth. (Hint: One megaton of TNT releases 4.2×1015 joules of energy.) Express your answer using two significant figures.
Answer #1

Given that

total mass of comet m = 4.2*1012 kg

speed of comet v = 2.5*104 m/s

(a)

total kinetic energy KE = 1/2*m*v2

KE = 0.5*4.2*1012*(2.5*104)2

KE = 13.12*1020 J

(b)

Given that

One megaton of TNT releases 4.2×1015 joules of energy. so,

total kinetic energy in TNT = 13.12*1020/4.2*1015

total kinetic energy in TNT = 312000 megaton of TNT